3(5^-x/4) = 15
a.-4
b.4
c.-1/3
d.1/3
Help, please? Show all steps.

Answer by ewatrrr(24785)


Hi
3(5^-x/4) = 15
5^-x/4 = 5
5^-x/4 = 5^1
-x/4 = 1
x = -4